Assume that the sales of a certain appliance dealer are approximated by a linear function. Suppose that sales were $6500 in 1982 and $64,500 in 1987. Let x =0 represent 1982. Find the equation giving yearly sales S(x).
Available choices:
S(x)= 11,600x + 6,500
S(x)= 58,000x + 64,500
S(x) = 58,000x + 6,500
S(x)= 11,600x + 64,500

Answer by gonzo(654)

answer is 11,600*x + 6,500.
-----
here's why.
1982 = 6500
1987 = 64500
total increase is 64500 - 6500 = 58000
total number of years from 1982 to 1987 is 1987 - 1982 = 5.
yearly increase in sales = total increase in sales divided by total increase in years.
this becomes 58000 / 5 = 11600.
since you start at 6500, the equation has to be 6500 + yearly increase of 11600.
equation would be y = 6500 + 11600 * x.
when x = 0, y = 6500 + 0 * 11600 = 6500 + 0 = 6500 which is the same as the value in 1982.
when x = 5, y = 6500 + 5 * 11600 = 6500 + 58000 = 64500 which is the same as the value in 1987.
-----
answer is 11,600 * x + 6,500.